OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = –9x + 4? A)(10, –86) B)(–4, –58) C)(6, –41) D)(–6, 57)

OpenStudy (anonymous):

-9(10) + 4 = -90 + 4 = -86 So A , (10,-86 is correct.